# TR-LLM: Integrating Trajectory Data for Scene-Aware LLM-Based Human Action Prediction

### v2

@@ I INTRODUCTION
Predicting human behavior is essential for AI systems to integrate seamlessly into our lives and provide effective support. This capability is particularly crucial for applications like support robots, which need to anticipate human actions and proactively perform tasks in a home environment. While traditional approaches have relied heavily on video data for action prediction, they often fall short due to their limited temporal and spatial scope. Human behavior is influenced by complex environmental and personal factors that extend beyond what can be captured in video footage. These include temporal elements like time of day, environmental conditions such as room temperature, the state and position of objects in the scene, and human-specific attributes including personal characteristics, action history, and the relationship with others present in the room. These factors exert significant influence on human behavior as exemplified by the contrast between retrieving an item from a fridge during a time-constrained morning routine versus during a more leisurely evening meal preparation or the contrast between weekday vs weekend routines. To truly understand and predict human behavior, we must take into account this broader contextual landscape, moving beyond the constraints of conventional video-based methods.
@@
Our work addresses the challenge by leveraging the capabilities of Large Language Models (LLMs), which possess broad knowledge of human behaviors and actions in home environments, and encodes the ability to plausibly predict human actions by considering diverse scene contexts, even in unseen scenarios.
However, observable scene contexts in the real world are often insufficient for predicting a person’s intentions accurately. For instance, scene context derived from sensor data (e.g., cameras, microphones) frequently suffers from limitations in coverage, sensitivity, or resolution, leading to increased uncertainty in predictions. This presents an inherent limitation in LLM-based action predictions.
To tackle the issue, we propose a multi-modal prediction framework that integrates physical constraints with LLM-based action prediction. Specifically, we use a person’s past trajectory to infer their likely destination, which imposes physical constraints on the person’s next target object, helping to narrow down potential actions (Fig.1). Since the target area prediction is intricately influenced by factors such as room layout and the person’s location and speed, we leverage an indoor human locomotion dataset to learn these relationships, allowing us to reliably derive a probabilistic distribution of target areas. To summarize our approach, we introduced a novel multi-modal action prediction method that integrates physical and semantic spaces. The key idea is to observe human actions from two spatially distinct perspectives—physical and semantic factors—which compensates for the limitations of each perspective and effectively refine the prediction. We believe that our multi-modal approach unlocks new potential for human action prediction in real-world applications.
@@
(i) Incorporating LLMs to human action anticipation considering a wide variety of semantic scene context
(ii) Proposing a multi-modal action prediction framework incorporating LLMs with trajectory that imposes physical constraints on the person’s next target object, helping to narrow down potential actions.
(iii) Building an evaluation dataset for our multi-modal approach, which includes scene map, trajectory, and scene contexts.
(iv) Demonstrating that the integration of LLM and trajectory data synergistically enhances performance, with the improvement being particularly pronounced when the LLM has limited scene information as input.
@@ II Related works
II-A Vision-based action prediction
Previous works have primarily focused on vision-based human action prediction. One line of research has focused on video-based action prediction, utilizing the recent past sequence of video frames to forecast future actions. Given the availability of large-scale action video datasets (e.g., Ego4D [2], Human3.6M [9], Home Action Genome [1]), many researchers have adopted machine-learning-based approaches [8, 7, 5, 4], often employing sequential models such as LSTM [6] or transformers [15, 14]. Some studies have integrated multimodal models to improve performance, combining video with other modalities (e.g., video-acoustic models [3], video-LLM models [11, 10]). Others have focused on object-based cues within the video to predict future actions [12, 13]. Another research direction involves predicting future motions from past motions, using only the human body’s pose data. For instance, [19] introduced a stochastic model for predicting diverse future actions with probability, while [17, 18] incorporated graph convolutional networks with attention mechanisms to improve prediction accuracy. Additionally, [16] proposed the use of a diffusion model to account for multi-person interactions.
In existing video-based action prediction research, a limitation lies in the spatial and temporal restrictions of scene context. Most studies focus on information within the limited field of view of ego-centric or third-person cameras, which only provides localized data. This narrow perspective hinders understanding of broader, more complex behaviors in real-world environments. Additionally, privacy concerns restrict the observation of individuals’ daily activities over extended periods in uncontrolled settings, limiting the ability to capture a diverse range of daily action patterns from existing datasets. To overcome these constraints, we propose leveraging large language models (LLMs), which encode vast everyday knowledge in natural language. By integrating LLMs, we remove spatial and temporal limitations, enabling a more flexible and generalized approach to action prediction that captures broader patterns applicable to diverse real-world scenarios.
II-B LLM-based action prediction
@@
To the best of our knowledge, only two studies have explored LLM-based human action prediction. [20] utilized an LLM to predict the next human activity by identifying the object in the environment that the person is most likely to interact with. [21] focused on LLM-based human action prediction to enable a robot to plan corresponding supportive actions. In their approach, they provided the LLM with object layouts and action history to predict the next human action. However, their problem settings are relatively simple, relying on limited cues for predicting the next target object. Moreover, while they account for scene context, they do not consider physical constraints such as human motion or the spatial structure of the environment.
@@ III-A Problem definition
In daily life, individuals perform sequential actions while interacting with various objects. For instance, one might take a glass from a cupboard, fill it with a drink from the fridge, and then sit on the couch to drink. In this study, we focus on predicting the transition between these sequential actions, specifically forecasting the next action as a person moves toward the target object. We assume the scene is a room-scale home environment with two persons performing separate actions, where we can observe various scene states except for the persons’ internal states (Fig.2).
@@ III-C1 LLM-based prediction
Fig.3 shows the pipeline of LLM-based target object prediction. In the initial step, various scene contexts—such as time of day, the person’s state (including location, action history, and conversation), and a list of objects—are provided as a text prompt and input into the LLM. Additionally, an order prompt is used to guide the LLM in generating a comprehensive, free-form list of potential target objects and their associated actions. This intermediate step facilitates the identification of probable target objects by considering their corresponding actions. Consequently, we derive a ranking of objects in the environment based on the target object candidates given in the previous step and the initial scene context prompt. The ranking is divided into four probability levels: A (high probability), B (moderate probability), C (low probability), and D (very low probability). The output from the LLM is subsequently converted into numerical scores and normalized into probabilities that sum to 1. In this study, scores of 15, 10, 5, and 1 were assigned to categories A, B, C, and D, respectively.
@@ III-C2 Trajectory-based prediction
We utilize past trajectory data and scene map information to constrain the target area of a person. Since the target area prediction is heavily influenced by factors such as room layout, the person’s location, and speed, we leverage the LocoVR dataset[32] to learn these intricate relationships. The LocoVR dataset consists of over 7,000 human trajectories captured in 131 indoor environments, with each trajectory segment spanning from a randomly assigned start to a goal location within the room. We employed a simple U-Net model to learn how to specify the possible target area based on the past trajectory and scene map (see Appendix-B2). Once the model predicts the probabilistic distribution of the target area, we calculate the overlap between this predicted distribution and the spatial location of objects in the scene to estimate the probability of each object being the target.
Fig.4 illustrates the process of trajectory-based prediction. In an early stage of trajectory (left image), the potential target area is broadly distributed. As the person progresses, the predicted target area and potential target objects are progressively narrowed down (center and right images). This refinement is based on a learned policy reflecting typical human behavior: when moving toward a specific goal, individuals generally avoid retracing their steps or taking unnecessary detours.
Figure 4: Trajectory-based goal prediction
@@ IV-A Evaluation data
Furthermore, we manually designed daily action scenarios, detailing factors such as day/time, persona, current location, target object, subsequent actions, conversations, and more. These scenarios were transformed into text prompts for input to the LLM. We generated a trajectory from the current location to the target object by leveraging a model trained on the LocoVR dataset[32], which incorporates the scene map to guide the path (see Appendix-C2 for details).
@@
IV-B Experimental setup
@@
We employed ChatGPT-4 to evaluate prediction performance using three types of input data to assess the model’s robustness under degraded conditions: (1) full scene context, (2) scene context without conversation, and (3) scene context without conversation and past action history. These variations simulate real-world scenarios where certain information is unavailable. The evaluation focuses on how our method mitigates this degradation to improve performance.
IV-C Evaluation metric
We conducted evaluations on both target object prediction and action prediction tasks.
Target object prediction: The goal of this task is to identify the correct target object from a set of 30-40 objects in each scene. In this evaluation, we considered the prediction correct if the groundtruth object appeared within the top three predicted objects.
Action prediction: This task involves predicting the action based on the target object and the scene contexts. We evaluated the similarity between the predicted and groundtruth actions, classifying them as correct or incorrect (see Appendix-B1 for details). Note that if the target object prediction was incorrect, the action prediction was also deemed incorrect, as it relied on the predicted target object.
@@ IV-C Evaluation metric
For both tasks, we calculated accuracy as the percentage of correct predictions across all scenarios. To ensure reproducibility, we conducted three trials for each evaluation and used the average of these trials as the final result.
IV-D Results
IV-D1 Quantitative result
Fig.6(a) illustrates the accuracy of target object prediction. While the accuracy of LLM-based predictions decreases as input scene information is limited, our method significantly mitigates this degradation. Furthermore, it is noteworthy that our method outperforms both the standalone performance of the LLM and the trajectory-based predictions, demonstrating a synergistic improvement in performance through the integration of LLM and trajectory data. Fig.6(b) presents the accuracy of action prediction based on the predicted target object. A similar trend is observed as in the target object prediction. Note that overall accuracy is lower than that of target object prediction because it is a subsequent process dependent on the target object prediction.
Fig.7(a) and 7(b) show the performance of target object prediction and action prediction as a function of the observed past trajectory length. The results demonstrate that performance improves for both tasks with longer observed trajectories. This is attributed to the increasing availability of cues for refining the target area as the trajectory progresses.
@@ IV-D1 Quantitative result
(a) Accuracy of target object prediction
(b) Accuracy of action prediction
Figure 6: Effect of our method for reducing input to the LLM: Our method mitigates the accuracy degradation of LLM-based predictions with reduced input, outperforming both LLM and trajectory-based predictions through the integration of LLM and trajectory data.
(a) Accuracy of target object prediction
(b) Accuracy of action prediction
Figure 7: Effect of our method for the length of trajectory: The performance improves for both tasks as the length of the observed trajectory increases. As the person moves, the predicted target area and potential target objects are progressively narrowed down.
IV-D2 Qualitative result
In this section, we present the qualitative performance of target object prediction using the LLM, trajectory, and our method through two sample scenarios.
LLM-based target object prediction: Fig.8 illustrates the results of target object prediction by the LLM, where the probabilities of each object are visualized on a map. Scenario 1 involves filling water at the sink to water a plant, while scenario 2 involves get a dirty glass at the table and take it to the sink. The results demonstrate that the objects relevant to the next action in each scenario are appropriately highlighted, despite a certain degree of uncertainty remaining in the predictions. This uncertainty arises from the limitation of predicting a person’s intention based on scene context from a semantic perspective.
Trajectory-based target object prediction: Fig.9 shows the results of target object prediction using trajectory data. The yellow distribution represents the predicted target area based on the observed trajectory, with each object color-coded from blue to red according to the degree of overlap with the predicted area (blue: low probability, red: high probability). In both scenario 1 and 2, the predicted target area initially covers a wide range shortly after the person starts from the sink. However, as the trajectory progresses, the target area becomes more focused, highlighting objects near the groundtruth. While the prediction roughly narrows down the target object region, it remains a significant challenge to precisely identify the groundtruth target object.
Target object prediction with our method: Fig.10 presents the results of target object prediction using our proposed method. In contrast to predictions using either the LLM or trajectory data alone, which retain some uncertainty, our method demonstrates a faster and more accurate narrowing of the target object. This indicates that by leveraging the complementary perspectives of the LLM’s semantic understanding and the trajectory’s physical context, the groundtruth target object is more effectively identified.
Figure 8: LLM-based target object prediction
Figure 9: Trajectory-based target object prediction
Figure 10: Trajectory+LLM-based target object prediction

### v3

@@ I INTRODUCTION
Predicting human behavior is essential for AI systems to integrate seamlessly into our lives and provide effective support. This capability is particularly crucial for applications like support robots, which need to anticipate human actions and proactively perform tasks in a home environment. While traditional approaches have relied heavily on video data for action prediction [6, 16], they often fall short due to their limited temporal and spatial scope. Human behavior is influenced by complex environmental and personal factors that extend beyond what can be captured in video footage. These include temporal elements like time of day, environmental conditions such as room temperature, the state and position of objects in the scene, and human-specific attributes including personal characteristics, action history, and the relationship with others present in the room. These factors exert significant influence on human behavior as exemplified by the contrast between retrieving an item from a fridge during a time-constrained morning routine versus during a more leisurely evening meal preparation or the contrast between weekday vs weekend routines. To truly understand and predict human behavior, we must take into account this broader contextual landscape, moving beyond the constraints of conventional video-based methods.
@@
Recent studies have explored the use of Large Language Models (LLMs) for predicting human actions [20, 21]. These models leverage extensive knowledge of human behavior in home environments to forecast actions across diverse scene contexts and generally perform robustly even in zero-shot settings. However, observable scene contexts in the real world are often insufficient for predicting a person’s intentions accurately. For instance, scene context derived from sensor data (e.g., cameras, microphones) frequently suffers from limitations in coverage, sensitivity, or resolution, leading to increased uncertainty in predictions. This presents an inherent limitation in LLM-based action predictions.
To address this shortcoming, we propose a multi-modal prediction framework that enhances LLM-based action prediction by integrating physical constraints. Our approach uses a person’s past trajectory to infer their likely destination, thereby imposing physical constraints on their next target object and narrowing down potential actions (Fig.1). Recognizing that target area prediction is highly dependent on factors such as room layout, location, and speed, we leverage an indoor human locomotion dataset [33] to learn these complex relationships, enabling us to derive a reliable probabilistic distribution of target areas.
Unlike vision-language models (VLMs) [34], which are adept at recognizing and interpreting images and text yet often struggle to predict physical quantities under complex image constraints, our method incorporates a dedicated physically-aware prediction module trained on practical datasets to overcome these challenges.
@@
  • Input: Incorporating LLMs to human action anticipation considering a wide variety of semantic scene context
  • Proposing a multi-modal action prediction framework incorporating LLMs with trajectory that imposes physical constraints on the person’s next target object, helping to narrow down potential actions.
  • Building an evaluation dataset for our multi-modal approach, which includes scene map, trajectory, and scene contexts.
  • Demonstrating that our method markedly enhances prediction performance compared with LLM and VLM, especially in scenarios where the LLM has limited access to scene information.
@@ II Related works
II-A Vision-based approach
Previous works have primarily focused on vision-based human action prediction. One line of research has focused on video-based action prediction, utilizing the recent past sequence of video frames to forecast future actions. Given the availability of large-scale action video datasets (e.g., Ego4D [2], Human3.6M [9], Home Action Genome [1]), many researchers have adopted machine-learning-based approaches [8, 7, 5, 4], often employing sequential models such as LSTM [6] or transformers [15, 14]. Some studies have integrated multimodal models to improve performance, combining video with other modalities (e.g., video-acoustic models [3], video-text models [11, 10]). Others have focused on object-based cues within the video to predict future actions [12, 13]. Another research direction involves predicting future motions from past motions, using only the human body’s pose data. For instance, [19] introduced a stochastic model for predicting diverse future actions with probability, while [17, 18] incorporated graph convolutional networks with attention mechanisms to improve prediction accuracy. Additionally, [16] proposed the use of a diffusion model to account for multi-person interactions.
In existing video-based action prediction research, a limitation lies in the spatial and temporal restrictions of scene context. Most studies focus on information within the limited field of view of ego-centric or third-person cameras, which only provides localized data. This narrow perspective hinders understanding of broader, more complex behaviors in real-world environments. Additionally, privacy concerns restrict the observation of individuals’ daily activities over extended periods in uncontrolled settings, limiting the ability to learn a diverse range of daily action patterns from existing datasets. To overcome these constraints, we propose leveraging large language models (LLMs), which encode vast everyday knowledge in natural language. By incorporating LLMs, we remove spatial and temporal limitations, enabling a more flexible and generalized approach to action prediction that captures broader patterns applicable to diverse real-world scenarios.
II-B Language model-based approach
@@
To the best of our knowledge, only a few studies have explored LLM-based human action prediction. [20] utilized an LLM to predict the next human activity by identifying the object in the environment that the person is most likely to interact with. [21] focused on LLM-based human action prediction to enable a robot to plan corresponding supportive actions. In their approach, they provided the LLM with object lists and action history to predict the next human action. However, while they account for scene context, they do not consider physical constraints such as human motion or the spatial structure of the environment.
Furthermore, vision-language models (VLMs)[34] have recently emerged as a prominent technology. These models possess the capability to recognize and understand both images and text, and they are utilized in a wide range of applications. Although they could potentially be applied to our task, pre-trained VLMs generally lack the ability to predict physical quantities under complex image constraints since they are primarily trained to establish correspondences between images and text. In contrast, our method can evaluate complex image constraints and predict the target object by employing a physically-aware prediction module that operates independently of the LLM module.
@@ III-A Problem definition
In daily life, individuals perform sequential actions while interacting with various objects. For instance, one might take a glass from a cupboard, fill it with a drink from the fridge, and then sit on the couch to drink. In this study, we focus on predicting the transition between these sequential actions, specifically forecasting the next action as a person moves toward the target object. We assume the scene is a room-scale home environment where we can observe various scene states except for the persons’ internal states (Fig.2). The input comprises both text-based and image-based scene contexts. The text-based context includes semantic details—such as the time of day, object lists, a person’s action history, and conversations—while the image-based context provides physical scene information through a 2D image that shows an object layout map and the person’s trajectory. In our task, we first predict the target object using both contexts and then forecast the future action based on this prediction.
@@ III-C1 LLM-based prediction
Fig.3 shows the pipeline of LLM-based target object prediction. In the initial step, various scene contexts—such as time of day, the person’s state (including location, action history, and conversation), and a list of objects in the room—are provided as a text prompt and input into the LLM. Additionally, an order prompt is used to guide the LLM in generating a comprehensive, free-form list of potential target objects and their associated actions. This intermediate step facilitates the identification of probable target objects by considering their corresponding actions. Consequently, we derive a ranking of objects in the environment based on the target object candidates given in the previous step and the initial scene context prompt. The ranking is divided into four probability levels: A (high probability), B (moderate probability), C (low probability), and D (very low probability). The output from the LLM is subsequently converted into numerical scores and normalized into probabilities that sum to 1. In this study, scores of 15, 10, 5, and 1 were assigned to categories A, B, C, and D, respectively.
@@ III-C2 Trajectory-based prediction
We utilize past trajectory data and scene map information to constrain the target area of a person. Since the target area prediction is heavily influenced by factors such as room layout, the person’s location, and speed, we leverage the LocoVR dataset[33] to learn these intricate relationships. The LocoVR dataset consists of over 7,000 human trajectories captured in 131 indoor environments, with each trajectory segment spanning from a randomly assigned start to a goal location within the room. We employed a simple U-Net model to identify potential target areas based on past trajectories and scene maps. Specifically, the model takes as input a binary obstacle scene map (HxWx1) and a trajectory represented as points on time-series images (HxWxt), and it outputs the target object’s position as an image (HxWx1). During training, we compute the binary cross entropy loss between the predicted and ground truth target object positions represented as the image. The trained model predicts the probabilistic distribution of the target area, and we calculate the overlap between this predicted distribution and the spatial location of objects in the scene to estimate the probability of each object being the target.
Fig.4 illustrates the process of trajectory-based target object prediction. In an early stage of trajectory (left image), the potential target area is broadly distributed. As the person progresses, the predicted target area and potential target objects are progressively narrowed down (center and right images). This refinement is based on a learned policy reflecting typical human behavior: when moving toward a specific goal, individuals generally avoid retracing their steps or taking unnecessary detours.
Figure 4: Trajectory-based target object prediction: At the beginning of the trajectory (left image), the potential target area is widely distributed. As the person advances, both the predicted target area and the range of potential target objects become progressively more focused (center and right images).
@@ IV-A Evaluation data
Furthermore, we manually designed daily action scenarios, detailing factors such as day/time, persona, current location, target object, subsequent actions, conversations, and more. These scenarios were transformed into text prompts for input to the LLM. We generated a trajectory from the current location to the target object by leveraging a model trained on the LocoVR dataset, which incorporates the scene map to guide the path (visit our website[32] for the detail).
@@
IV-B Benchmarks
Since action prediction using room-scale scene context is a novel task, few existing studies could directly apply to our work. Therefore, we introduce the following three baseline methods as benchmarks:
LLM: In this method, the LLM predicts the target object and its corresponding human action based on a text-based scene context. The input comprises day/time information, action history, conversation, and an object candidate list. The LLM produces scores for each target object and selects the predicted action associated with the highest-scoring target object. This approach is equivalent to the LLM-based action prediction modules described in [20, 21], though our implementation differs because the original codes were unavailable.
VLM: We employed pre-trained VLM models to benchmark baseline performance on multimodal data. The VLM input consists of a text-based scene context and an image that displays the scene map and the trajectory. In the image, each object and trajectory is uniquely color-coded, and a color list of the objects (e.g., sink: cyan, table: gold, trajectory: red) is included in the text-based scene context to help the model accurately identify object locations and trajectories. The VLM then outputs scores for the target objects and selects the corresponding action for the most plausible object, similar to an LLM.
Trajectory: We evaluated a standalone trajectory-based method, which is a component of our approach. This serves as an ablation study to assess how our method improves performance on the target object prediction compared with standalone trajectory-based method.
For both LLM and VLM, we used two types of pre-trained models—ChatGPT-4o-mini and Llama3.2—to evaluate performance across different model architectures.
IV-C Experimental setup
@@
We provided three types of text-based scene context as input data to assess the model’s robustness under degraded conditions: (1) full scene context, (2) scene context without conversation, and (3) scene context without conversation and past action history. These variations simulate real-world scenarios where certain information may be unavailable. The evaluation examines how our method mitigates performance degradation, demonstrating its robustness to noise in practical settings. Additionally, we evaluate performance under three trajectory progress conditions: d¿1m, d¿2m, and d¿3m. The performance differences across these conditions reveal how the trajectory information contributes in each method.
IV-D Evaluation tasks
@@
We conducted evaluations on both target object prediction and action prediction tasks. For both tasks, we conducted three trials for each evaluation and used the average of these trials as the final result to ensure the reproducibility.
Target object prediction: This task serves as an essential intermediate module for action prediction. The objective is to identify the ground truth target object from a set of 30–40 objects in each scene. For evaluation, we used the top-5 accuracy metric, which considers a prediction correct if the ground truth object appears among the top five predicted objects.
Action prediction: In this task, we evaluate action prediction accuracy by computing the cosine similarity between the predicted actions and their corresponding ground truth values using the SBERT library.
IV-E Results
IV-E1 Quantitative result
@@
TableI presents the accuracy of target object prediction. We describe the following conclusions based on the tableI:
(a) Our method outperforms both LLM and VLM across GPT and Llama models. Notably, as the available text-based scene context decreases, the prediction accuracy of both LLM and VLM deteriorates significantly, whereas our method maintains robust performance. This demonstrates that the physical cues extracted from trajectories effectively compensate for the reduced semantic context.
(b) VLM performance is comparable to—or slightly better than—that of LLM, and both degrade similarly as the available scene context decreases. These results suggest that pre-trained VLMs may struggle with solving physical problems under complex geometric constraints, which could limit their added value when combined with text-based scene contexts. This is because pre-trained VLMs are not designed to predict physical quantities under complex image constraints, as they are primarily trained to establish correspondences between images and text.
(c) Notably, our method’s performance significantly exceeds the average of the standalone LLM and Trajectory approaches, indicating that integrating LLM and trajectory information effectively compensates for individual weaknesses while leveraging their respective strengths.
(d) Regarding trajectory progress distance (d), performance improves as d increases because longer trajectories offer more spatial cues to narrow down potential destination areas as shown in Fig4. This trend is evident in both our method and the standalone trajectory approach. In contrast, VLM does not exhibit a clear correlation between traveled distance and accuracy, suggesting that it struggles to process physical information effectively.
TableII displays the performance on action prediction. A similar trend to that observed in target object prediction is evident, emphasizing that the target object is a key cue for forecasting future actions. These results underscore the effectiveness of our approach compared to other baselines.
TABLE I: Accuracy of target object prediction
$Method$ $Accuracy[\%]$
All wo/conv wo/conv-hist
LLM-llama 43.2 35.0 26.5
VLM-llama ($d>1m$) 40.3 31.6 31.4
VLM-llama ($d>2m$) 40.9 32.2 31.2
VLM-llama ($d>3m$) 41.4 32.0 31.2
Ours ($d>1m$)
(LLM-llama+Trajectory) 47.9 45.9 41.4
Ours ($d>2m$)
(LLM-llama+Trajectory) 49.9 48.1 43.5
Ours ($d>3m$)
(LLM-llama+Trajectory) 51.4 49.8 46.2
LLM-gpt 44.7 37.2 28.0
VLM-gpt ($d>1m$) 50.9 40.9 27.8
VLM-gpt ($d>2m$) 51.5 42.4 30.3
VLM-gpt ($d>3m$) 51.6 43.0 31.0
Ours ($d>1m$)
(LLM-gpt+Trajectory) 54.6 51.2 44.3
Ours ($d>2m$)
(LLM-gpt+Trajectory) 56.4 53.5 47.3
Ours ($d>3m$)
(LLM-gpt+Trajectory) 58.6 56.3 49.1
Trajectory ($d>1m$) 33.3 33.3 33.3
Trajectory ($d>2m$) 36.7 36.7 36.7
Trajectory ($d>3m$) 39.7 39.7 39.7
TABLE II: Accuracy of action prediction
$Method$ $Cosinesimularity$
All wo/conv wo/conv-hist
LLM-llama 0.220 0.201 0.176
VLM-llama ($d>1m$) 0.208 0.199 0.176
VLM-llama ($d>2m$) 0.207 0.198 0.187
VLM-llama ($d>3m$) 0.203 0.200 0.185
Ours ($d>1m$)
(LLM-llama+Trajectory) 0.218 0.210 0.207
Ours ($d>2m$)
(LLM-llama+Trajectory) 0.222 0.214 0.210
Ours ($d>3m$)
(LLM-llama+Trajectory) 0.227 0.218 0.213
LLM-gpt 0.335 0.278 0.236
VLM-gpt ($d>1m$) 0.368 0.313 0.256
VLM-gpt ($d>2m$) 0.370 0.311 0.254
VLM-gpt ($d>3m$) 0.370 0.305 0.243
Ours ($d>1m$)
(LLM-gpt+Trajectory) 0.387 0.353 0.310
Ours ($d>2m$)
(LLM-gpt+Trajectory) 0.397 0.365 0.322
Ours ($d>3m$)
(LLM-gpt+Trajectory) 0.409 0.379 0.331
IV-E2 Qualitative result
In this section, we first present visual comparisons between our method and baseline approaches in sample scenes (Fig.6), and then detail how target object prediction evolves as the trajectory progresses (Fig.7).
Fig.6 displays the target object prediction results using LLM-gpt, VLM-gpt, standalone trajectory-based method (Trajectory), and our proposed method (LLM-gpt+Trajectory). The red point indicates the starting position of the trajectory, while the green line and point represent the observed trajectory and current position, respectively. The yellow distribution illustrates the predicted target area based on the observed trajectory, while the objects—color-coded from blue to red—indicate predicted target object probabilities from low to high. The figure indicates that the LLM assigns high probabilities to several objects, including the target; however, some mispredictions persist due to the inherent difficulty of inferring a person’s intentions solely from the semantic scene context. Similarly, the VLM produces some incorrect predictions because of its limited capability in physically-aware prediction. Although the exact mechanism remains unclear for the pre-trained model, we observe that the VLM likely to assign relatively high probabilities to objects near the trajectory, indicating that it does not appear to predict the future target area. On the other hand, the standalone trajectory-based target object prediction assigns high probabilities widely around the ground truth target area, yet it struggles to accurately pinpoint the target object—especially when the trajectory has just begun. By leveraging both semantic and physical cues, our method more effectively narrows down the target object compared to using either LLM or trajectory data alone.
Fig.7 illustrates the detailed behavior of the trajectory-based target object prediction by showing two distinct trajectory patterns over time within the same room layout. Initially (upper row), the target area probability distributions (shown in yellow) are broadly spread in both scenarios. As the trajectories progress (middle and bottom rows), these distributions narrow to regions near the ground truth target object, demonstrating the successful refinement process over time.
Figure 6: Target object prediction: LLM and VLM identify several object candidates, but their predictions include errors due to limited physically-aware capabilities. In contrast, while the standalone trajectory-based method roughly predicts target object areas, it lacks precision. Our method narrows down the target object more effectively by leveraging both semantic and physical cues.
Figure 7: Behavior of trajectory-based target object prediction over time: In the top row, the target area probability distributions (highlighted in yellow) are widely spread across both scenarios. As the trajectory progresses, the distributions in the middle and bottom rows become increasingly concentrated around the ground truth target object areas.